# Water Evolution & Inventories of Super-Earths Orbiting Late M-Dwarfs

### v1

@@ 2 Methods
The model used in this study is a modified version of that in Moore, Cowan, & Boukaré (2023), without a persistent basal magma ocean; we instead test various magma ocean water partition coefficients across different planetary masses, and a basal magma ocean would be represented by a higher partition coefficient. Moore, Cowan, & Boukaré (2023) showed that a short-lived basal magma ocean could benefit surface habitability, which could help rapidly-desiccated planets if substantial water becomes sequestered in the mantle; however, a long-lived basal magma ocean leads to the majority of water becoming trapped in the mantle. Basal magma oceans are probably more likely for super-Earths (C.-É Boukaré, priv. comm.), which we discuss in Section 4.
@@ 2.1 Expected Mass Dependence
Mass Habitable Zone (HZ) Distance [AU] Period [days] EL during RG DL rest of sim Total Maximum
1 $M_{\oplus}$ Hot, Inner Edge 0.025 4.81 7.72 2.82 10.5 19.7
@@
Table 1: Expected amount of water that will be lost as a function of planetary mass, $M_{\mathrm{p}}$, and orbital distance, neglecting a magma ocean and deep-water cycle. Each HZ orbital distance fixed throughout a given simulation, and determined using the calculator of Kopparapu et al. (2013) at 4.5 Gyr, comparable to the age of the Solar System; the Inner Edge corresponds to the runaway greenhouse limit, the Outer Edge corresponds to freeze-out of carbon dioxide, and the Middle is taken to be the average orbital distance between these two limits. Here, “EL” refers to energy-limited loss, which is expected to decrease with $M_{\mathrm{p}}$; “RG” is the runaway greenhouse; and “DL” is diffusion-limited loss, expected to increase with $M_{\mathrm{p}}$. The total expected loss is the sum of the previous two columns, while the maximum expected loss is calculated using the maximum of EL and DL loss at each timestep. In practice, these planets lose much less water due to the cycling of water between interior, surface, and atmosphere.
@@ 4 Discussion
In this study we have adopted the HZ limits of Kopparapu et al. (2013), but habitable planets orbit very close to their M-dwarf hosts and are likely tidally-locked (e.g., Kasting et al. 1993). GCM simulations suggest that the inner edge moves inwards to 2$\times$ the stellar flux at the runaway greenhouse limit (Yang et al., 2013, 2014). We therefore perform a sensitivity analysis by instead taking the inner edge to be at 2$\times$ the RG limit of Kopparapu et al. (2013), i.e., $a_{\mathrm{InnerHZ,tl}}=\sqrt{(a_{\mathrm{InnerHZ}}^{2})/2}=0.017$ AU. For this closer Inner HZ limit, our model predicts that the planet is permanently in a runaway greenhouse and never becomes habitable (cf. Turbet et al. 2023). In principle, a wider habitable zone incorporating a closer Inner HZ edge could benefit the super-Earths in our results (see Fig. 5) if the planet is able to exit the runaway greenhouse phase while retaining a water inventory.
@@
However, $D$ only controls partitioning during the relatively short-lived magma ocean phase; less of the total water inventory may be lost to space than expected, but water sequestered in the mantle does not necessarily enhance surface habitability. All else remaining the same, the amount of water present within the solid mantle by 5 Gyr is nearly identical regardless of $D$ due to the short lifetime of the magma ocean compared to the much longer deep-water cycling phase and the significant degassing from the hot mantle following surface solidification, so varying $D$ has nearly no effect on the final planetary water inventories, their surface water regime, nor their habitability prospects. A Gyr-lived basal magma ocean — posited by Moore, Cowan, & Boukaré (2023) for an Earth-mass planet, and probably more likely to exist in super-Earths (C. É. Boukaré, priv. comm.) — would presumably cause even more water to become trapped in the solid mantle than our current simulations indicate. This reinforces our conclusion that substantial water becomes locked in the solid mantle for super-Earths, in agreement with Schaefer & Sasselov (2015).
@@ 5 Summary & Conclusions
Fig. 4, instead initializing all planets with 2–8 Earth Oceans, reveals that surface desiccation is more likely for super-Earths due to the enhanced diffusion-limited loss during deep-water cycling, but also because more water becomes sequestered in the solid mantle. In this dry, in-situ formation scenario (e.g., Schneeberger et al. 2024), super-Earths are more likely to become desiccated and less likely to end up in an Earth-like surface water regime.
@@ Acknowledgments
KM and NBC thank Sean Raymond and Charles-Édouard Boukaré for useful exchanges, and the anonymous referees for insightful comments and suggestions. NBC acknowledges support from an NSERC Discovery Grant, a Tier 2 Canada Research Chair, and an Arthur B. McDonald Fellowship. The authors also thank the Trottier Space Institute (TSI) and l’Institut de recherche sur les exoplanètes (iREx) for their financial support and dynamic intellectual environment. This research was partially funded by the Heising-Simons Foundation.

### v2

# Water Evolution & Inventories of Super-Earths Orbiting Late M Dwarfs

@@ 2 Methods
The model used in this study is a modified version of that in Moore, Cowan, & Boukaré (2023), without a persistent basal magma ocean; we instead test various magma ocean water partition coefficients across different planetary masses, and a basal magma ocean would be represented by a higher partition coefficient. Moore, Cowan, & Boukaré (2023) showed that a short-lived basal magma ocean could benefit surface habitability, which could help rapidly-desiccated planets if substantial water becomes sequestered in the mantle; however, a long-lived basal magma ocean leads to the majority of water becoming trapped in the mantle. Basal magma oceans are probably more likely for super-Earths (C.-É Boukaré, private communication, 2024), which we discuss in Section 4.
@@ 2.1 Expected Mass Dependence
Mass [$M_{\oplus}$] Habitable Zone (HZ) Distance [AU] Period [days] EL during RG DL rest of sim Total Maximum
1 Hot, Inner Edge 0.025 4.81 7.72 2.82 10.5 19.7
@@
Table 1: Expected amount of water that will be lost as a function of planetary mass, $M_{\mathrm{p}}$, and orbital distance, neglecting a magma ocean and deep-water cycle. Each HZ orbital distance fixed throughout a given simulation, and determined using the calculator of Kopparapu et al. (2013) at 4.5 Gyr, comparable to the age of the Solar System; the Inner Edge corresponds to the runaway greenhouse limit, the Outer Edge corresponds to freeze-out of carbon dioxide, and the Middle is taken to be the average orbital distance between these two limits. Here, “EL” refers to energy-limited loss, which is expected to decrease with $M_{\mathrm{p}}$; “RG” is the runaway greenhouse; “DL” is diffusion-limited loss, expected to increase with $M_{\mathrm{p}}$; and “Rest of Sim” refers to the remainder of the simulation, following the end of the magma ocean/runaway greenhouse period. The total expected loss is the sum of the previous two columns, while the maximum expected loss is calculated using the maximum of EL and DL loss at each timestep. In practice, these planets lose much less water due to the cycling of water between interior, surface, and atmosphere.
@@ 4 Discussion
In this study we have adopted the HZ limits of Kopparapu et al. (2013), but habitable planets orbit very close to their M-dwarf hosts and are likely tidally-locked (e.g., Kasting et al. 1993). Global climate model (GCM) simulations suggest that the inner edge moves inwards to 2$\times$ the stellar flux at the runaway greenhouse limit (Yang et al., 2013, 2014). We therefore perform a sensitivity analysis by instead taking the inner edge to be at 2$\times$ the RG limit of Kopparapu et al. (2013), i.e., $a_{\mathrm{InnerHZ,tl}}=\sqrt{(a_{\mathrm{InnerHZ}}^{2})/2}=0.017$ AU. For this closer Inner HZ limit, our model predicts that the planet is permanently in a runaway greenhouse and never becomes habitable (cf. Turbet et al. 2023). In principle, a wider habitable zone incorporating a closer Inner HZ edge could benefit the super-Earths in our results (see Fig. 5) if the planet is able to exit the runaway greenhouse phase while retaining a water inventory.
@@
However, $D$ only controls partitioning during the relatively short-lived magma ocean phase; less of the total water inventory may be lost to space than expected, but water sequestered in the mantle does not necessarily enhance surface habitability. All else remaining the same, the amount of water present within the solid mantle by 5 Gyr is nearly identical regardless of $D$ due to the short lifetime of the magma ocean compared to the much longer deep-water cycling phase and the significant degassing from the hot mantle following surface solidification, so varying $D$ has nearly no effect on the final planetary water inventories, their surface water regime, nor their habitability prospects. ”A basal magma ocean that persists for Gyr-timescales — posited by Moore, Cowan, & Boukaré (2023) for an Earth-mass planet, and probably more likely to exist in super-Earths (C. É. Boukaré, private communication, 2024) — would presumably cause even more water to become trapped in the solid mantle than our current simulations indicate. This reinforces our conclusion that substantial water becomes locked in the solid mantle for super-Earths, in agreement with Schaefer & Sasselov (2015).
@@ 5 Summary & Conclusions
Fig. 4, which instead shows results for all planets initialized with 2–8 Earth Oceans, independent of planetary mass, reveals that surface desiccation is more likely for super-Earths due to the enhanced diffusion-limited loss during deep-water cycling, but also because more water becomes sequestered in the solid mantle. In this dry, in-situ formation scenario (e.g., Schneeberger et al. 2024), super-Earths are more likely to become desiccated and less likely to end up in an Earth-like surface water regime.
@@ Acknowledgments
K. M. and N. B. C. thank Sean Raymond and Charles-Édouard Boukaré for useful exchanges, and the anonymous referees for insightful comments and suggestions. K. M. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) Postgraduate Scholarships-Doctoral (PGS D) Fellowship. N. B. C. acknowledges support from an NSERC Discovery Grant, a Tier 2 Canada Research Chair, and an Arthur B. McDonald Fellowship. The authors also thank the Trottier Space Institute (TSI) and l’Institut de recherche sur les exoplanètes (iREx) for their financial support and dynamic intellectual environment. This research was partially funded by the Heising-Simons Foundation.